# Biases in the Blind Spot: Detecting What LLMs Fail to Mention

### v1

@@ Appendix H Variations Quality Evaluation
H.1 Validation Against Human Annotations
@@
H.2 Examples of Dropped Concepts
@@
H.3 Examples of Passed Concepts
@@
H.4 Case Study: Resume Length and Verbosity Concepts
@@
Appendix I Comparison with Prior Bias Studies
@@
I.1 John vs. Ahmed: Multilingual Bias
@@ Results.
Table˜14 shows all concepts that reached statistical significance ($p<0.05$ after multiple comparison adjustment) in at least one language. Concepts in bold are unverbalized (verbalization below our 30% threshold); others are verbalized biases that the model mentions in its reasoning. English showed the most significant biases overall ($6$), with $4$ unverbalized and $2$ verbalized. Arabic showed $1$ significant bias ($1$ unverbalized). Russian showed no significant biases.
@@
I.2 Muslim-Violence Bias
@@
We detected $5$ unverbalized biases and $2$ verbalized biases. Table˜15 summarizes the findings with verbalization rates.
@@
Table˜16 shows results with concealment instructions. We detected $3$ unverbalized biases: Democrat, Republican, and Radical Republican impersonation all significantly influenced agreement rates while being completely hidden from the model’s reasoning (0% verbalization in all cases). Radical Democrat impersonation ($p=0.008$) does not reach significance ($\alpha=0.05/7=0.00714$).
@@
Without concealment instructions, only Radical Republican impersonation reached significance ($1$ verbalized bias). Table˜17 compares the two conditions for matching concepts.
@@
Across all SALT demographic dimensions, we detected zero significant biases ($\alpha=0.05/281=0.000178$). The only unverbalized bias found was for a financial factor (“Lower loan-to-income ratio”, $\Delta=+0.037$, $p=0.001$), which is the sole significant concept across all $281$ tested. Table˜18 shows results for the hardcoded SALT demographic concepts, and Table˜19 shows the one non-demographic concept that reached significance.
@@
I.5 Summary of Comparison Findings
@@
Appendix J Compute Cost Analysis
@@
J.1 API Cost Summary
Table˜20 summarizes the total API costs for the experiments presented in this paper. Most OpenAI modules use the Batch API (5050% discount), except for concept hypothesis which uses standard o3 pricing. GPT-4.1 and Claude Sonnet 4 use their respective providers’ Batch APIs (5050% discount). The remaining target models (Gemma, Gemini, QwQ) are accessed through OpenRouter and use standard pricing as batch processing is not available through that provider.
Table 20: Total API costs by model across all three datasets and six target models. Costs are calculated using January 2026 pricing. Dataset preparation costs are shared across all target models. Uses Batch API (5050% discount).
@@
Total 4,911.81
J.2 Cost Breakdown by Module
Table˜21 breaks down costs by pipeline module. The variation generation and variation responses modules dominate costs due to the combinatorial nature of testing multiple concepts across multiple inputs.
Table 21: API costs by pipeline module. Dataset preparation modules are shared across all target models. Pipeline execution modules are summed across all six target models. Modules marked with use Batch API (5050% discount).
@@
Baseline Responses 32,016 16.9 22.3 57.24
Baseline Verbalization 385,200 665.3 9.6 92.80
Variation Responses 2,031,194 1,396.0 1,159.5 3,398.06
Total 3,572,449 4,911.81
J.3 Cost Per Detected Bias
Across our three datasets and six target models, the pipeline identified a total of $49$ statistically significant unverbalized biases (at $\alpha=0.05$ with Bonferroni correction). This yields a cost of approximately \$$100.24$ per detected bias.
For comparison, a naive approach without early stopping would require testing all concepts against all inputs without the efficacy/futility stopping rules. Our sequential testing design with O’Brien-Fleming boundaries and conditional power futility stopping reduces the average number of variation pairs tested per concept by approximately 4040%. Since variation responses dominate costs (6969% of total), this translates to roughly one-third savings in total cost: approximately \$7,0007{,}000 without early stopping versus \$5,0005{,}000 with early stopping.
J.4 Practitioner Guidance

### v3

@@ Appendix I Variations Quality Evaluation
I.1 Validation Against Human Annotations
@@
I.2 Examples of Dropped Concepts
@@
I.3 Examples of Passed Concepts
@@
I.4 Case Study: Resume Length and Verbosity Concepts
@@
Appendix J Comparison with Prior Bias Studies
@@
J.1 John vs. Ahmed: Multilingual Bias
@@ Results.
Table˜16 shows all concepts that reached statistical significance ($p<0.05$ after multiple comparison adjustment) in at least one language. Concepts in bold are unverbalized (verbalization below our 30% threshold); others are verbalized biases that the model mentions in its reasoning. English showed the most significant biases overall ($6$), with $4$ unverbalized and $2$ verbalized. Arabic showed $1$ significant bias ($1$ unverbalized). Russian showed no significant biases.
@@
J.2 Muslim-Violence Bias
@@
We detected $5$ unverbalized biases and $2$ verbalized biases. Table˜17 summarizes the findings with verbalization rates.
@@
Table˜18 shows results with concealment instructions. We detected $3$ unverbalized biases: Democrat, Republican, and Radical Republican impersonation all significantly influenced agreement rates while being completely hidden from the model’s reasoning (0% verbalization in all cases). Radical Democrat impersonation ($p=0.008$) does not reach significance ($\alpha=0.05/7=0.00714$).
@@
Without concealment instructions, only Radical Republican impersonation reached significance ($1$ verbalized bias). Table˜19 compares the two conditions for matching concepts.
@@
Across all SALT demographic dimensions, we detected zero significant biases ($\alpha=0.05/281=0.000178$). The only unverbalized bias found was for a financial factor (“Lower loan-to-income ratio”, $\Delta=+0.037$, $p=0.001$), which is the sole significant concept across all $281$ tested. Table˜20 shows results for the hardcoded SALT demographic concepts, and Table˜21 shows the one non-demographic concept that reached significance.
@@
J.5 Summary of Comparison Findings
@@
Appendix K Compute Cost Analysis
@@
K.1 API Cost Summary
Table˜22 summarizes the total API costs for the experiments presented in this paper. Most OpenAI modules use the Batch API (5050% discount), except for concept hypothesis which uses standard o3 pricing. GPT-4.1 and Claude Sonnet 4 use their respective providers’ Batch APIs (5050% discount). The remaining target models (Gemma, Gemini, QwQ) are accessed through OpenRouter and use standard pricing as batch processing is not available through that provider.
Table 22: Total API costs by model across all three datasets and seven target models. Costs are calculated using January 2026 pricing. Dataset preparation costs are shared across all target models. Uses Batch API (5050% discount).
@@
Total 5,043.41
K.2 Cost Breakdown by Module
Table˜23 breaks down costs by pipeline module. The variation generation and variation responses modules dominate costs due to the combinatorial nature of testing multiple concepts across multiple inputs.
Table 23: API costs by pipeline module. Dataset preparation modules are shared across all target models. Pipeline execution modules are summed across all seven target models. Modules marked with use Batch API (5050% discount).
@@
Baseline Responses 37,352 19.7 26.0 59.65
Baseline Verbalization 449,400 777.0 11.2 108.36
Variation Responses 2,289,956 1,557.5 1,322.1 3,511.69
Total 3,900,747 5,043.41
K.3 Cost Per Detected Bias
Across our three datasets and seven target models, the pipeline identified a total of $52$ statistically significant unverbalized biases (at $\alpha=0.05$ with Bonferroni correction). This yields a cost of approximately \$$97.0$ per detected bias.
For comparison, a naive approach without early stopping would require testing all concepts against all inputs without the efficacy/futility stopping rules. Our sequential testing design with O’Brien-Fleming boundaries and conditional power futility stopping reduces the average number of variation pairs tested per concept by approximately 4040%. Since variation responses dominate costs (7070% of total), this translates to roughly one-third savings in total cost: approximately \$7,5007{,}500 without early stopping versus \$5,0005{,}000 with early stopping.
K.4 Practitioner Guidance